Prove that (R^5)/R ~_ R^4 as rings
1
Expert's answer
2020-04-16T18:46:33-0400

Let's consider the following structure of R5, R4 and R:


R5 = {(r1, r2, r3, r4, r5) | ri ∈ R for i = 1,2,3,4,5},

R4 = {(r1, r2, r3, r4) | ri ∈ R for i= 1,2,3,4},

R = {(0, 0, 0, 0, r) | r ∈ R }, with ring operations applied to each component(i.e. (r1, r2, r3, r4, r5) +(or *) (p1, p2, p3, p4, p5) = (r1+(*)p1, ..., r5+(*)p5) and same for R4 and R).


It could be inferred(from the expressions, represented above) that R5/R = {(r1, r2, r3, r4, 0) | ri ∈ R for i = 1, 2, 3, 4}. Here is the isomorphism between the R5/R and R4:


f(r1, r2, r3, r4, 0) = (r1, r2, r3, r4).

f-1(r1, r2, r3, r4) = (r1, r2, r3, r4, 0).

Let's check f and f-1 are homomorphisms of rings:


Since the last component of the element from R5/R always equals to zero(and due to our definition of ring operations) f is a homomorphism:

f(a + b) = f(a) + f(b)

f(a*b) = f(a)*f(b) where a,b belong to the R5/R


f-1 = g is just adding the 0 as the last componet. So, because of the definition of ring operations =>

g(a + b) = g(a) + g(b)

g(a*b) = g(a)*g(b) where a,b belong to the R4


This means, that f is invertible homomorphism => R5/R ~f R4.